OpenStudy (anonymous):

Solve the inequality and write your answer in interval notation. 9x+16 greater than or equal to -9+6x

OpenStudy (anonymous):

\[9x+16 \ge -9+6x\] Combine like terms: \[9x - 6x \ge -9-16\] \[= 3x \ge -25\] Divide by 3: \[= x \ge -25/3\] And now in interval notation: \[x: [-25/3,\infty)\]